Ticket Number: HEA-NOC/20060728-1 Ticket Status: CLOSE
Ticket Type: scheduled Resolver: HEA-NOC
Ticket Opened: 20060728 09:42 UTC+1 Problem Start: 20060728 11:00 UTC+1
Ticket Closed: 20060821 17:18 UTC+1 Problem End: 20060803 14:00 UTC+1
The circuit to NUI is being upgraded to 10 Mbps. This involves:
- Install a new circuit
- Provision a new client premises HEAnet router
- Test circuit
- Change routing to use a static route over new connection
- Cease old circuit
20060728 09:36 BB BT report new circuit is installed. Scheduled time with
NUI and Kerna (NUI's IT support) to install new router (11am, 28/7/2006).
20060731 12:05 BB BT have repaired a config problem with the new circuit,
and it is now passing IP traffic. The NUI CPE router is not properly
authenticating users - debugging.
20060802 11:05 BB Circuit tested for 24 hours, and looks good. Scheduling
routing changeover with Kerna for Thursday 3/8/2006 at 13:00.
20060803 09:56 BB Upgrade confirmed with client for 13:00.
20060804 10:16 BB Routing changes complete.
20060821 17:17 BB Traffic routing normally, closing ticket.
This ticket can be monitored at http://www.hea.net/tickets/20060728-1